OpenStudy (anonymous):

multiply and write the answer without exponents (2x10^3)x(3x10^6)

OpenStudy (raden):

we can multiply 2 by 3, first. it is 6 then 10^3 = 1,000 and 10^6 = 1,000,000. multiply of them

OpenStudy (anonymous):

so 6,000,000

OpenStudy (raden):

6 x 1,000, x 1,000,000 (look how many zeroes are there, 9 right ?) = 6,000,000,000

OpenStudy (anonymous):

ok so 2x3=6 and 10^3+10^6= 10^9

OpenStudy (raden):

10^3 multiply 10^6, right ?

OpenStudy (raden):

if u use the property of exponent : a^m * a^n = a^(m+n)

OpenStudy (raden):

so, 10^3 * 10^6 = 10^(3+6) = 10^9

OpenStudy (anonymous):

ok got it!!